Night Plan UT 31-Dec-2020



Split Night: 1st half LBTI, 2nd half D/E (this plan)

Night Center 00:22

Handover 02:17

Morning Astronomical Twilight: 05:54

Non-photometric all night

Link to Google Docs here

Night Plan

  • DONE Re-Deploy TCS 2021A after handover from LBTI to LBTO D/E
  • SOUL AO Checkout following problems of 20201220 UT- CRITICAL - up to 2hours
    • DONE OSCO at handover - 30min - NCPA on / NCPA off
    • AO Scipts: /home/lbto/Service_Observing/D_E_Time_UT_Dec28_Jan_02/AO_Testing/
    • DONE Bin3 Test 0.5-0.75 hr - HIGH priority
      • closed loop data 6 positons of the instrument rotator
      • test the camera lens limits with new M3 position
      • Can use either IDL presets, or AO stars from OT in above directory
    • Bin 1-4 tests on-sky - check functionality
      • use scripts generated from OT in above directory
    • Other Issues:
      • IT 8282 - Bin? pupil files issue
      • IT 8281 - AGW1 loses star at DETXY 10" offset
      • IT 8284 - GOPT issues (pending daytime work too)
      • IT 8271 - preset gains for ESM needed - work on creating the necessary values
    • Observe ESM fields using full AO but with N375 camera - 1 hour per field
      • M67
      • HCG 31
    • Additional?
  • ESM checkout - (M3 and AO checkout must be completed first) - if not started or completed roll over to another night
    • Scripts: /home/lbto/Service_Observing/D_E_Time_UT_Dec28_Jan_02/ESM_Testing
    • on-sky checks Bin 2 400Hz (FLAO config) vs Bin 2 100Hz (SOUL config) - 1-2 hrs
    • Collect ESM data - <1" seeing - stable - 1hr minmum
      • Palomar 2 - if clear take photometric standards - probing sensitivity vs FLAO - sets at 3am
  • MODS Checkout - Medium Priority
    • Verify Rotator ZP with MOS masks (20min)
    • Fields:
      • StoneA - sets 10pm
      • Stone D1f - midnight-2am
      • Stone E1f - midnight-5am
      • Stone I - 3am-twilight
      • MMS files in mms/ sub-directory
    • Collect verification AGW transform (1 hour)
    • Check NS dithers
      • In /home/lbto/Service_Observing/D_E_Time_UT_Dec28_Jan_02/MODS_NSdithers, there are 3 scripts for end-of-night target, Semele (V<~15), available from 2am to morning:
        • async.txt - grid of small async dithers
        • sync+async - large sync dither followed by an updatepointing and a grid of detxy abs dithers
        • sync+async_rel large sync dither followed by a grid of detxy rel dithers.
  • Poor Weather Backups
    • ESM data collection: this is different than ESM tests above - this is sytematic data colleciton - above is checking on paramters for optimal use
      • Scripts: /home/lbto/Service_Observing/D_E_Time_UT_Dec28_Jan_02/ESM_Testing
      • DONE M67 (3 filters and bright and faint AO Ref star) - K-band only completed
      • DONE NGC 3921 all filters
    • ICRF3 Spectroscopy (Rothberg)
      • Scripts: /home/lbto/Service_Observing/Nov_Dec_2020_E_D_Nights/Poor_Weather/ICRF3_Rothberg_2020B
      • ICR5 93.24-13.51 Highest Priority - cycles through G210 K,H,J, and z (fraternal or mono)
        • sets at 3am
        • Target can sustain 2mags of extinction or more - uses 0.25" slit - J=12, H=9.5, K=7.2
      • Observe telluric first
      • Double Check high Dec objects to make sure telluric/science are close spatially
    • If clear but poor seeing, please observe standard stars, cycling through z,J,H,K,Ks
      • Scripts: /home/lbto/Service_Observing/Nov_Dec_2020_E_D_Nights/Poor_Weather/LUCI_Phot_ZP
      • Select stars to use based on visibility - OT file is in directory, export the scripts as needed
      • Please observe the same star at least two times at different airmasses
-- BarryRothberg - 30 Dec 2020
Topic revision: r6 - 01 Jan 2021, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback